# Highest-weight vectors and three-point functions in GKO coset decomposition

### v1

@@ 6.2 Integral
Let $N\in\mathbb{Z}_{\geq 0}$, $l,m,n\in\frac{1}{2}\mathbb{Z}_{\geq 0}$ such that $l+m+n\in\mathbb{Z}$ and $m+l+n\leq N$. We want to compute integral
@@
The Selberg type integral (6.19) equals
@@
$\langle u_{m}(\mu)|\mathcal{U}_{\nu,n}(1,1)|u_{l}(\lambda)\rangle\\ =\sum_{r\geq 0}\binom{\nu+2n}{r}\langle v_{m}(\mu),e^{g_{0}}\Big{(}e^{-g_{0}}e^{n\sqrt{2}\phi(1)}(-1)^{nh_{0}^{(1)}(h_{0}^{(1)}-1)}e^{g_{0}}\mathcal{O}^{(N)}_{r;\nu}(1)\Big{)}e^{-g_{0}}v_{l}(\lambda)\rangle\\ =\sum_{r\geq 0}\binom{\nu+2n}{r}\int\limits_{0\leq t_{N}\leq\dots\leq t_{1}\leq 1}\!\!\!\!\langle v_{m}(\mu),\Big{(}\gamma^{r}(1)e^{n\sqrt{2}\phi(1)}(-1)^{nh_{0}^{(1)}(h_{0}^{(1)}-1)}e^{\frac{\nu}{\sqrt{2\kappa}}\varphi(1)}\Big{)}\\ \prod_{i=1}^{N}e^{g_{0}}S(t_{i})e^{-g_{0}}\,v_{l}(\lambda)\rangle\prod_{i=1}^{N}dt_{i}$ (6.21)
Recall that $S(t)=\beta(t)e^{-\frac{2}{\sqrt{2\kappa}}\varphi(t)}$. Then $e^{g_{0}}S(t)e^{-g_{0}}=(\beta(t)-e^{\sqrt{2}\phi(t)})e^{-\frac{2}{\sqrt{2\kappa}}\varphi(t)}$.
$\!\!\!\!\!\!\int\limits_{0\leq t_{N}\leq\dots\leq t_{1}\leq 1}\!\!\!\!\!\!\!\!\!\!\langle v_{m}(\mu),\Big{(}\gamma^{r}(1)e^{n\sqrt{2}\phi(1)}(-1)^{nh_{0}^{(1)}(h_{0}^{(1)}{-}1)}e^{\frac{\nu}{\sqrt{2\kappa}}\varphi(1)}\Big{)}\prod_{i=1}^{N}(\beta(t_{i})-e^{\sqrt{2}\phi(t_{i})})e^{-\frac{2}{\sqrt{2\kappa}}\varphi(t_{i})}\,v_{l}(\lambda)\rangle\prod_{i=1}^{N}dt_{i}\\ =(-1)^{N-r+4n(m-n)(m-n-\frac{1}{2})}\sum_{I\subset\{1,\dots,N\}}\;\;\int\limits_{0\leq t_{N}\leq\dots\leq t_{1}\leq 1}\langle v_{m}(\mu),\gamma^{r}(1)e^{n\sqrt{2}\phi(1)}e^{\frac{\nu}{\sqrt{2\kappa}}\varphi(1)}\\ \times\prod_{i=1}^{N}\Big{(}(1-\delta_{i\in I})\beta(t_{i})+\delta_{i\in I}e^{\sqrt{2}\phi(t_{i})}\Big{)}e^{-\frac{2}{\sqrt{2\kappa}}\varphi(t_{i})}\,v_{l}(\lambda)\rangle\prod_{i=1}^{N}dt_{i}.$ (6.22)
@@
$\langle v_{m}(\mu),\gamma^{r}(1)e^{n\sqrt{2}\phi(1)}e^{\frac{\nu}{\sqrt{2\kappa}}\varphi(1)}\prod_{i=1}^{N}\Big{(}(1-\delta_{i\in I})\beta(t_{i})+\delta_{i\in I}e^{\sqrt{2}\phi(t_{i})}\Big{)}e^{-\frac{2}{\sqrt{2\kappa}}\varphi(t_{i})}\,v_{l}(\lambda)\rangle\\ =(-1)^{r}r!\prod_{i=1}^{N}t_{i}^{-\frac{\lambda}{\kappa}}(1-t_{i})^{-\frac{\nu}{\kappa}-1}\prod_{1\leq i<j\leq N}(t_{i}-t_{j})^{\frac{2}{\kappa}}\prod_{i\in I}t_{i}^{2l}(1-t_{i})^{2n+1}\prod_{i,j\in I,\,i<j}(t_{i}-t_{j})^{2}.$ (6.23)
Now we can replace $(t_{i}-t_{j})^{\frac{2}{\kappa}}$ by $|t_{i}-t_{j}|^{\frac{2}{\kappa}}$ and the integral domain by cube $[0,1]^{N}$ with additional factor $N!$. Then the value of the integral does not depend on the choice of $I$ and we can assume that $I=\{1,\dots,N-r\}$ with additional factor $\binom{N}{r}$. Finally we get integral (6.19), namely
$\frac{\tilde{C}_{m,n,l}(\mu,\nu,\lambda)}{\|\tilde{u}_{l}(\lambda)\|^{2}\|\tilde{u}_{n}(\nu)\|^{2}}=\langle u_{m}(\mu)|\mathcal{U}_{\nu,n}(1,1)|u_{l}(\lambda)\rangle\\ =(-1)^{N+4n(m-n)(m-n-\frac{1}{2})}\binom{\nu+2n}{N-m+n+l}\frac{1}{(m-n-l)!}S^{N}_{-n,m,-l}\left(-\frac{\lambda}{\kappa}+1,-\frac{\nu}{\kappa},\frac{1}{\kappa}\right).$ (6.24)
@@ 6.3 Symmetries, constant term, particular cases
$G_{N,R,r}(\alpha,g)=\left[\begin{aligned} &\prod\limits_{j=0}^{N-1}\!\Gamma\big{(}\alpha{+}jg{+}\mathds{1}_{j>N-R}(j{-}(N-R)){+}\mathds{1}_{j<2r}(j{-}2r)\big{)}&\text{ if $R>0$}\\ &\frac{\prod\limits_{j=0}^{N-R-1}\!\!\Gamma\big{(}\alpha{+}jg{+}\mathds{1}_{j<2r}(j{-}2r)\big{)}}{\prod\limits_{j=N}^{N-R-1}\Gamma\big{(}\alpha{+}jg{+}(j{-}(N{-}R))\big{)}}&\text{ if $R\leq 0$}\end{aligned}\right.$ (6.30)
$\tilde{G}_{N,R,r}(\alpha,g)=\left[\begin{aligned} &\frac{\prod\limits_{j=-R+1}^{-1}\!\!\Gamma\big{(}\alpha+jg+R+j\big{)}}{\prod\limits_{j=-R+1}^{N-1}\Gamma\big{(}\alpha+jg+\mathds{1}_{j>N-2r}(j+2r-N)\big{)}}&\text{ if $R>0$}\\ &\prod\limits_{j=0}^{N-1}\!\Gamma\big{(}\alpha+jg+\mathds{1}_{j<-R}(j+R)+\mathds{1}_{j>N-2r}(j+2r-N)\big{)}^{-1}&\text{ if $R\leq 0$}\end{aligned}\right.$ (6.31)
@@
$S^{N}_{n,m,l}(\alpha,\beta,g)=\Gamma(1+m+l+n)\prod_{j=0}^{N-1}\frac{\Gamma\big{(}1+(j+1)g+\mathds{1}_{j>N{-}m{-}l{-}n}(j{-}N{+}m{+}l{+}n)\big{)}}{\Gamma(1+g)}\\ G_{N,m+n-l,l}(\alpha,g)G_{N,l+m-n+1,n-1/2}(\beta,g)\tilde{G}_{N,m-n-l,m}(\alpha+\beta+(N-1)g,g)$ (6.32)
@@
It is straightforward to compare this with the formula 6.32.
@@
In the case $a^{\prime}=b^{\prime}=0$ the constant term (6.35) coincides with the one conjectured by Forrester [For95] and proven in [KNPV15, Th. 6.2] (more precisely on have to take $q\rightarrow 1$ limit of $m=0$ case of this theorem)
@@
In the case case $a^{\prime}=1,b^{\prime}=0$ the constant term coincides with another particular case of [KNPV15, Th. 6.2], namely one have to take $q\rightarrow 1$ limit and set $n_{0}=n-m$ in notations of loc. cit
@@ Appendix A Monodromy cancellation
Recall notations for hypergeometric function introduced in formula (4.34). Recall also transformation $\hat{I}$ defined in (4.35). In order to write monodromy of hypergeometric function we will need also transformations $\hat{R},\hat{S}$
$\hat{R}\vec{a}=(a_{1},a_{1}-a_{3}+1,a_{1}-a_{2}+1),\quad\hat{S}\vec{a}=\vec{a}^{\prime}=(a_{2},a_{1},a_{3}).$ (A.1)
@@
is rational function by $z$ with poles in $0,1$ and $\infty$, if and only if
@@
There is a correspondence between terms in right side of formulas (A.6) and (A.7), namely
${}_{2}F_{1}(\hat{R}\vec{a}|z^{-1}){}_{2}F_{1}(\hat{R}(-\vec{a}-\vec{r})|z^{-1})$ $={}_{2}F_{1}(\hat{R}\hat{I}\vec{a}|z^{-1}){}_{2}F_{1}(\hat{R}\hat{I}(-\vec{a}-\vec{r})|z^{-1})$ (A.8)
${}_{2}F_{1}(\hat{R}\hat{S}\vec{a}|z^{-1}){}_{2}F_{1}(\hat{R}\hat{S}(-\vec{a}-\vec{r})|z^{-1})$ $={}_{2}F_{1}(\hat{R}\hat{S}\hat{I}\vec{a}|z^{-1}){}_{2}F_{1}(\hat{R}\hat{S}\hat{I}(-\vec{a}-\vec{r})|z^{-1})$ (A.9)
${}_{2}F_{1}(\hat{R}\hat{S}\vec{a}|z^{-1}){}_{2}F_{1}(\hat{R}(-\vec{a}-\vec{r})|z^{-1})$ $={}_{2}F_{1}(\hat{R}\hat{S}\hat{I}\vec{a}|z^{-1}){}_{2}F_{1}(\hat{R}\hat{I}(-\vec{a}-\vec{r})|z^{-1})$ (A.10)
${}_{2}F_{1}(\hat{R}\vec{a}|z^{-1}){}_{2}F_{1}(\hat{R}\hat{S}(-\vec{a}-\vec{r})|z^{-1})$ $={}_{2}F_{1}(\hat{R}\hat{I}\vec{a}|z^{-1}){}_{2}F_{1}(\hat{R}\hat{S}\hat{I}(-\vec{a}-\vec{r})|z^{-1})$ (A.11)
@@
Due to our assumptions, the third and fourth terms in the right side of formulas (A.6) and (A.7) should cancel each other. The cancellation of the third term gives
@@
On the other hand, the function $H_{\vec{r}}(\vec{a}|z)$ can have singularities only at $z=0,1,\infty$ and these singularities are branching points. For $p_{\vec{r}}(\vec{a})$ given by formula (A.5) the arguments above shows that the function $H_{\vec{r}}(\vec{a}|z)$ at $z=0,\infty$ can have only poles. Hence the monodromy of $H_{\vec{r}}(\vec{a}|z)$ at $z=1$ is trivial. Therefore the singularities at $z=1$ are also just poles. Hence $H_{\vec{r}}(\vec{a}|z)$ is is a rational function. ∎
@@ Appendix B Three point functions
Let $\chi$ be a (probably infinite) sum of terms of the form $e^{-\tau\xi}$, where $\tau$ is a formal parameter. By conjugation $\chi^{*}$ we denote operation which acts as $e^{-\tau\xi}\mapsto e^{\tau\xi}$. Let us introduce the function
@@
$\mathsf{E}\left[\sum e^{\tau\xi_{i}}-\sum e^{\tau\eta_{j}}\right]=\frac{\prod\eta_{j}}{\prod\xi_{i}}.$ (B.2)
@@
$\mathsf{C}^{\mathrm{Vir}}(u_{1},u_{2},u_{3};q_{1},q_{2})=\mathsf{E}\left[\frac{q_{1}^{2}q_{2}^{2}(u_{1}u_{2}u_{3}-u_{2}^{2}-u_{3}^{2})+u_{1}u_{2}^{-1}u_{3}^{-1}+u_{1}^{-1}u_{2}u_{3}^{-1}+u_{1}^{-1}u_{2}^{-1}u_{3}-u_{1}^{-2}}{(1-q_{1})(1-q_{2})}\right]$ (B.3)
$\mathsf{C}^{\mathfrak{sl}}(u_{1},u_{2},u_{3};q_{1},q_{2})=\mathsf{E}\left[\frac{q_{1}^{2}q_{2}(u_{1}u_{2}u_{3}-u_{1}^{2}-u_{2}^{2})+u_{1}u_{2}^{-1}u_{3}^{-1}+u_{1}^{-1}u_{2}u_{3}^{-1}+u_{1}^{-1}u_{2}^{-1}u_{3}-u_{1}^{-2}}{(1-q_{1})(1-q_{2})}\right]$ (B.4)
@@
where $v=e^{\tau\alpha},q_{1}=e^{\tau\epsilon_{1}},q_{2}=e^{\tau\epsilon_{2}}$. Using this we have
@@
where in first transformation we used notations $u_{1}=e^{\tau a_{1}},u_{2}=e^{\tau a_{2}},u_{3}=e^{\tau a_{3}},q_{1}=e^{\tau\epsilon_{1}},q_{2}=e^{\tau\epsilon_{2}}$ and in second transformation we used Theorems 4.5, Corollary 3.17, and identification of parameters
@@
Using these functions we can renormalize four-point conformal blocks. For the algebra $\widehat{\mathfrak{sl}}(2)$ we define as follows
@@
where $u_{i}=e^{\tau\lambda_{i}}$, $u^{\prime}=e^{\tau\lambda^{\prime}}$, $q_{1}=e^{2\tau}$, $q_{2}=e^{-2\kappa\tau}$, c.f. (B.7). For the Virasoro case we defined
@@
$\mathsf{C}^{\mathfrak{sl}}(u_{1},u_{2},u_{3};q_{1},q_{2})=\mathsf{C}^{\mathfrak{sl}}(u_{1},u_{2},u_{3};q_{1},q_{1}^{-1}q_{2})\mathsf{C}^{\mathrm{Vir}}(u_{1},u_{2},u32;q_{1}q_{2}^{-1},q_{2}).$ (B.10)
The $q_{1},q_{2}$ parameters here exactly correspond to coset relations, e.g. (4.59), Namely, we take, $q_{1}=e^{2\tau}$, $q_{2}=e^{-2(k+2)\tau}$, then for $\mathsf{C}^{\mathfrak{sl}}$ on right side parameters are $e^{2\tau},e^{-2(k+3)\tau}$ i,e, correspond to $\widehat{\mathfrak{sl}}(2)_{k+1}$ and for $\mathsf{C}^{\mathrm{Vir}}$ parameters are $e^{2(k+3)\tau},e^{-2(k+2)\tau}$ i.e. correspond to Virasoro algebra with $b=b_{GKO}=-\mathrm{i}\sqrt{\frac{k+2}{k+3}}$. Relation on parameters $u$ also agrees with relation $P(\lambda)=P_{GKO}(\lambda)=-\frac{\lambda+1}{2(k+2)}b_{GKO}$.
@@
$\Psi_{k}^{\mathrm{full}}\left(\vec{\mu},\lambda;x,z\right)=\sum_{l\in\mathbb{Z}}\Psi_{k+1}^{\mathrm{full}}\left(\vec{\mu},\lambda+2l;x,z\right)\mathrm{F}_{b}^{\mathrm{full}}\left(\vec{P},P(\lambda)+lb;z\right),$ (B.11)

### v3

@@ 6.2 Integral
Let $N\in\mathbb{Z}_{\geq 0}$, $l,m,n\in\frac{1}{2}\mathbb{Z}_{\geq 0}$ such that $l+m+n\in\mathbb{Z}$ and $l+m+n\leq N$. We want to compute integral
@@
Note that $S^{N}_{0,0,0}(\alpha,\beta,g)$ is a standard Selberg integral, see e.g. [FW08].
@@
The Selberg-type integral (6.19) is equal to
@@
$\Big{\langle}u_{m}(\mu),\mathcal{U}_{\nu,n}(1,1)\,u_{l}(\lambda)\Big{\rangle}\\ =\sum_{r\geq 0}\binom{\nu+2n}{r}\Big{\langle}v_{m}(\mu),{\mathrm{e}}^{g_{0}}\Big{(}{\mathrm{e}}^{-g_{0}}:\!{\mathrm{e}}^{n\sqrt{2}\phi(1)}\!\!:(-1)^{nh_{0}^{(1)}(h_{0}^{(1)}-1)}{\mathrm{e}}^{g_{0}}\mathcal{O}^{(N)}_{r;\nu}(1)\Big{)}{\mathrm{e}}^{-g_{0}}v_{l}(\lambda)\Big{\rangle}\\ =\sum_{r\geq 0}\binom{\nu+2n}{r}\int\limits_{0\leq t_{N}\leq\dots\leq t_{1}\leq 1}\!\!\!\!\Big{\langle}v_{m}(\mu),\Big{(}\gamma^{r}(1):\!{\mathrm{e}}^{n\sqrt{2}\phi(1)}\!\!:(-1)^{nh_{0}^{(1)}(h_{0}^{(1)}-1)}:\!{\mathrm{e}}^{\frac{\nu}{\sqrt{2\kappa}}\varphi(1)}\!\!:\Big{)}\\ \prod_{i=1}^{N}{\mathrm{e}}^{g_{0}}S(t_{i}){\mathrm{e}}^{-g_{0}}\,v_{l}(\lambda)\Big{\rangle}\prod_{i=1}^{N}dt_{i}$ (6.21)
Recall that $\tilde{S}(t)={\mathrm{e}}^{g_{0}}S(t){\mathrm{e}}^{-g_{0}}=(\beta(t)-{\mathrm{e}}^{\sqrt{2}\phi(t)}):\!{\mathrm{e}}^{-\frac{2}{\sqrt{2\kappa}}\varphi(t)}\!\!:$, see formula (3.53). Then we can rewrite integral in the right side
$\!\!\!\!\!\!\int\limits_{0\leq t_{N}\leq\dots\leq t_{1}\leq 1}\!\!\!\!\!\!\!\!\!\!\Big{\langle}v_{m}(\mu),\Big{(}\gamma^{r}(1):\!{\mathrm{e}}^{n\sqrt{2}\phi(1)}\!\!:(-1)^{nh_{0}^{(1)}(h_{0}^{(1)}{-}1)}:\!{\mathrm{e}}^{\frac{\nu}{\sqrt{2\kappa}}\varphi(1)}\!\!:\Big{)}\\ \prod_{i=1}^{N}(\beta(t_{i})-:\!{\mathrm{e}}^{\sqrt{2}\phi(t_{i})}\!\!:):\!{\mathrm{e}}^{-\frac{2}{\sqrt{2\kappa}}\varphi(t_{i})}\!\!:v_{l}(\lambda)\Big{\rangle}\prod_{i=1}^{N}dt_{i}\\ =(-1)^{N-r+4n(m-n)(m-n-\frac{1}{2})}\sum_{I\subset\{1,\dots,N\}}\;\;\int\limits_{0\leq t_{N}\leq\dots\leq t_{1}\leq 1}\Big{\langle}v_{m}(\mu),\gamma^{r}(1):\!{\mathrm{e}}^{n\sqrt{2}\phi(1)}\!\!:\,:\!{\mathrm{e}}^{\frac{\nu}{\sqrt{2\kappa}}\varphi(1)}\!\!:\\ \times\prod_{i=1}^{N}\Big{(}(1-\delta_{i\in I})\beta(t_{i})+\delta_{i\in I}:\!{\mathrm{e}}^{\sqrt{2}\phi(t_{i})}\!\!:\Big{)}:\!{\mathrm{e}}^{-\frac{2}{\sqrt{2\kappa}}\varphi(t_{i})}\!\!:v_{l}(\lambda)\Big{\rangle}\prod_{i=1}^{N}dt_{i}.$ (6.22)
@@
$\Big{\langle}v_{m}(\mu),\gamma^{r}(1):\!{\mathrm{e}}^{n\sqrt{2}\phi(1)}\!\!:\,:\!{\mathrm{e}}^{\frac{\nu}{\sqrt{2\kappa}}\varphi(1)}\!\!:\prod_{i=1}^{N}\Big{(}(1-\delta_{i\in I})\beta(t_{i})+\delta_{i\in I}:\!{\mathrm{e}}^{\sqrt{2}\phi(t_{i})}\!\!:\Big{)}:\!{\mathrm{e}}^{-\frac{2}{\sqrt{2\kappa}}\varphi(t_{i})}\!\!:v_{l}(\lambda)\Big{\rangle}\\ =(-1)^{r}r!\prod_{i=1}^{N}t_{i}^{-\frac{\lambda}{\kappa}}(1-t_{i})^{-\frac{\nu}{\kappa}-1}\prod_{1\leq i<j\leq N}(t_{i}-t_{j})^{\frac{2}{\kappa}}\prod_{i\in I}t_{i}^{2l}(1-t_{i})^{2n+1}\prod_{i,j\in I,\,i<j}(t_{i}-t_{j})^{2}.$ (6.23)
Now we can replace $(t_{i}-t_{j})^{\frac{2}{\kappa}}$ by $|t_{i}-t_{j}|^{\frac{2}{\kappa}}$ and the integral domain by a cube $[0,1]^{N}$ with additional factor $N!$. Then the value of the integral does not depend on the choice of $I$ and we can assume that $I=\{1,\dots,N-r\}$ with additional factor $\binom{N}{r}$. Finally we get integral (6.19), namely
$\frac{\tilde{C}_{m,n,l}(\mu,\nu,\lambda)}{\|\tilde{u}_{l}(\lambda)\|^{2}\|\tilde{u}_{n}(\nu)\|^{2}}=\Big{\langle}u_{m}(\mu),\mathcal{U}_{\nu,n}(1,1)\,u_{l}(\lambda)\Big{\rangle}\\ =(-1)^{N+4n(m-n)(m-n-\frac{1}{2})}\binom{\nu+2n}{N-m+n+l}\frac{1}{(m-n-l)!}S^{N}_{-n,m,-l}\left(-\frac{\lambda}{\kappa}+1,-\frac{\nu}{\kappa},\frac{1}{\kappa}\right).$ (6.24)
@@ 6.3 Symmetries, constant term, particular cases
$G_{N,R,r}(\alpha,g)=\left[\begin{aligned} &\prod\limits_{j=0}^{N-1}\!\Gamma\big{(}\alpha{+}jg{+}\mathds{1}_{j>N-R}(j{-}(N-R)){+}\mathds{1}_{j<2r}(j{-}2r)\big{)}&\text{ if $R>0$},\\ &\frac{\prod\limits_{j=0}^{N-R-1}\!\!\Gamma\big{(}\alpha{+}jg{+}\mathds{1}_{j<2r}(j{-}2r)\big{)}}{\prod\limits_{j=N}^{N-R-1}\Gamma\big{(}\alpha{+}jg{+}(j{-}(N{-}R))\big{)}}&\text{ if $R\leq 0$},\end{aligned}\right.$ (6.30)
$\tilde{G}_{N,R,r}(\alpha,g)=\left[\begin{aligned} &\frac{\prod\limits_{j=-R+1}^{-1}\!\!\Gamma\big{(}\alpha+jg+R+j\big{)}}{\prod\limits_{j=-R+1}^{N-1}\Gamma\big{(}\alpha+jg+\mathds{1}_{j>N-2r}(j+2r-N)\big{)}}&\text{ if $R>0$},\\ &\prod\limits_{j=0}^{N-1}\!\Gamma\big{(}\alpha+jg+\mathds{1}_{j<-R}(j+R)+\mathds{1}_{j>N-2r}(j+2r-N)\big{)}^{-1}&\text{ if $R\leq 0$}.\end{aligned}\right.$ (6.31)
@@
$S^{N}_{n,m,l}(\alpha,\beta,g)=\Gamma(1+m+l+n)\prod_{j=0}^{N-1}\frac{\Gamma\big{(}1+(j+1)g+\mathds{1}_{j>N{-}m{-}l{-}n}(j{-}N{+}m{+}l{+}n)\big{)}}{\Gamma(1+g)}\\ G_{N,m+n-l,l}(\alpha,g)G_{N,l+m-n+1,n-1/2}(\beta,g)\tilde{G}_{N,m-n-l,m}(\alpha+\beta+(N-1)g,g).$ (6.32)
@@
It is straightforward to compare this with the formula (6.32).
@@
In the case $a^{\prime}=b^{\prime}=0$ the constant term (6.35) coincides with the one conjectured by Forrester [For95] and proven in [KNPV15, Th. 6.2] (more precisely one has to take $q\rightarrow 1$ limit of $m=0$ case of this theorem)
@@
In the case $a^{\prime}=1,b^{\prime}=0$ the constant term coincides with another particular case of [KNPV15, Th. 6.2], namely one has to take $q\rightarrow 1$ limit and set $n_{0}=n-m$ in notations of loc. cit
@@ Appendix A Monodromy cancellation
Recall notations for hypergeometric function introduced in formula (4.34). Recall also transformation $\hat{I}$ defined in (4.35). In order to write monodromy of hypergeometric function we will need also transformations $\hat{R}$ and $\hat{S}$
$\hat{R}\vec{a}=(a_{1},a_{1}-a_{3}+1,a_{1}-a_{2}+1),\quad\hat{S}\vec{a}=(a_{2},a_{1},a_{3}).$ (A.1)
@@
is a rational function of $z$ with poles in $0,1$ and $\infty$, if and only if
@@
There is a correspondence between terms in right sides of formulas (A.6) and (A.7), namely
${}_{2}F_{1}(\hat{R}\vec{a}|z^{-1}){}_{2}F_{1}(\hat{R}(-\vec{a}-\vec{r})|z^{-1})$ $={}_{2}F_{1}(\hat{R}\hat{I}\vec{a}|z^{-1}){}_{2}F_{1}(\hat{R}\hat{I}(-\vec{a}-\vec{r})|z^{-1}),$ (A.8)
${}_{2}F_{1}(\hat{R}\hat{S}\vec{a}|z^{-1}){}_{2}F_{1}(\hat{R}\hat{S}(-\vec{a}-\vec{r})|z^{-1})$ $={}_{2}F_{1}(\hat{R}\hat{S}\hat{I}\vec{a}|z^{-1}){}_{2}F_{1}(\hat{R}\hat{S}\hat{I}(-\vec{a}-\vec{r})|z^{-1}),$ (A.9)
${}_{2}F_{1}(\hat{R}\hat{S}\vec{a}|z^{-1}){}_{2}F_{1}(\hat{R}(-\vec{a}-\vec{r})|z^{-1})$ $={}_{2}F_{1}(\hat{R}\hat{S}\hat{I}\vec{a}|z^{-1}){}_{2}F_{1}(\hat{R}\hat{I}(-\vec{a}-\vec{r})|z^{-1}),$ (A.10)
${}_{2}F_{1}(\hat{R}\vec{a}|z^{-1}){}_{2}F_{1}(\hat{R}\hat{S}(-\vec{a}-\vec{r})|z^{-1})$ $={}_{2}F_{1}(\hat{R}\hat{I}\vec{a}|z^{-1}){}_{2}F_{1}(\hat{R}\hat{S}\hat{I}(-\vec{a}-\vec{r})|z^{-1}),$ (A.11)
@@
Due to our assumptions, the third and fourth terms in the right sides of formulas (A.6) and (A.7) should cancel each other. The cancellation of the third term gives
@@
On the other hand, the function $H_{\vec{r}}(\vec{a}|z)$ can have singularities only at $z=0,1,\infty$ and these singularities are branching points. For $p_{\vec{r}}(\vec{a})$ given by formula (A.5) the arguments above shows that the function $H_{\vec{r}}(\vec{a}|z)$ at $z=0,\infty$ can have only poles. Hence the monodromy of $H_{\vec{r}}(\vec{a}|z)$ at $z=1$ is trivial. Therefore the singularities at $z=1$ are also just poles. Hence $H_{\vec{r}}(\vec{a}|z)$ is a rational function. ∎
@@ Appendix B Three point functions
Let $\chi$ be a (probably infinite) sum of terms of the form ${\mathrm{e}}^{-\tau\xi}$, where $\tau$ is a formal parameter. By conjugation $\chi^{*}$ we denote operation which acts as ${\mathrm{e}}^{-\tau\xi}\mapsto{\mathrm{e}}^{\tau\xi}$. Let us introduce the function (plethystic exponent)
@@
$\mathsf{E}\left[\sum{\mathrm{e}}^{\tau\xi_{i}}-\sum{\mathrm{e}}^{\tau\eta_{j}}\right]=\frac{\prod\eta_{j}}{\prod\xi_{i}}.$ (B.2)
@@
$\mathsf{C}^{\mathrm{Vir}}(u_{1},u_{2},u_{3};q_{1},q_{2})=\mathsf{E}\left[\frac{q_{1}^{2}q_{2}^{2}(u_{1}u_{2}u_{3}-u_{2}^{2}-u_{3}^{2})+u_{1}u_{2}^{-1}u_{3}^{-1}+u_{1}^{-1}u_{2}u_{3}^{-1}+u_{1}^{-1}u_{2}^{-1}u_{3}-u_{1}^{-2}}{(1-q_{1})(1-q_{2})},\right]$ (B.3)
$\mathsf{C}^{\mathfrak{sl}}(u_{1},u_{2},u_{3};q_{1},q_{2})=\mathsf{E}\left[\frac{q_{1}^{2}q_{2}(u_{1}u_{2}u_{3}-u_{1}^{2}-u_{2}^{2})+u_{1}u_{2}^{-1}u_{3}^{-1}+u_{1}^{-1}u_{2}u_{3}^{-1}+u_{1}^{-1}u_{2}^{-1}u_{3}-u_{1}^{-2}}{(1-q_{1})(1-q_{2})}\right].$ (B.4)
@@
where $v={\mathrm{e}}^{\tau\alpha},q_{1}={\mathrm{e}}^{\tau\epsilon_{1}},q_{2}={\mathrm{e}}^{\tau\epsilon_{2}}$. Using this we have
@@
where in first transformation we used notations $u_{1}={\mathrm{e}}^{\tau a_{1}},u_{2}={\mathrm{e}}^{\tau a_{2}},u_{3}={\mathrm{e}}^{\tau a_{3}},q_{1}={\mathrm{e}}^{\tau\epsilon_{1}},q_{2}={\mathrm{e}}^{\tau\epsilon_{2}}$ and in second transformation we used Theorems 4.5, Corollary 3.18, and identification of parameters
@@
Using these functions we can renormalize four-point conformal blocks. For the algebra $\widehat{\mathfrak{sl}}(2)$ we define it as follows
@@
where $u_{i}={\mathrm{e}}^{\tau\lambda_{i}}$, $u^{\prime}={\mathrm{e}}^{\tau\lambda^{\prime}}$, $q_{1}={\mathrm{e}}^{2\tau}$, $q_{2}={\mathrm{e}}^{-2\kappa\tau}$, c.f. (B.7). For the Virasoro case we define
@@
$\mathsf{C}^{\mathfrak{sl}}(u_{1},u_{2},u_{3};q_{1},q_{2})=\mathsf{C}^{\mathfrak{sl}}(u_{1},u_{2},u_{3};q_{1},q_{1}^{-1}q_{2})\mathsf{C}^{\mathrm{Vir}}(u_{1},u_{2},u_{3};q_{1}q_{2}^{-1},q_{2}).$ (B.10)
The $q_{1},q_{2}$ parameters here exactly correspond to coset relations, e.g. (4.59), Namely, we take, $q_{1}={\mathrm{e}}^{2\tau}$, $q_{2}={\mathrm{e}}^{-2(k+2)\tau}$, then for $\mathsf{C}^{\mathfrak{sl}}$ on the right side parameters are ${\mathrm{e}}^{2\tau},{\mathrm{e}}^{-2(k+3)\tau}$ i.e. correspond to $\widehat{\mathfrak{sl}}(2)_{k+1}$ and for $\mathsf{C}^{\mathrm{Vir}}$ parameters are ${\mathrm{e}}^{2(k+3)\tau},{\mathrm{e}}^{-2(k+2)\tau}$ i.e. correspond to Virasoro algebra with $b=b_{GKO}=-{\mathrm{i}}\sqrt{\frac{k+2}{k+3}}$. Relation on parameters $u$ also agrees with relation $P(\lambda)=P_{GKO}(\lambda)=-\frac{\lambda+1}{2(k+2)}b_{GKO}$.
@@
$\Psi_{k}^{\mathrm{full}}\left(\vec{\mu},\lambda;x,z\right)=\sum_{l\in\mathbb{Z}}\Psi_{k+1}^{\mathrm{full}}\left(\vec{\mu},\lambda+2l;x,z\right)\mathrm{F}_{b}^{\mathrm{full}}\left(\vec{P},P(\lambda)+lb;z\right).$ (B.11)
@@
Data Availability
The authors declare that the data supporting the findings of this study are available within the paper.
Conflict of Interest
The authors have no relevant financial or non-financial interests to disclose.